Make x the subject of the following equations
x/b+x/b=1



Answer by Theo(13342)

equation is x/b + x/b = 1
since x/b + x/b = 2x/b, your equation becomes 2x/b = 1
multiply both side by b to get 2x = b
divide both sides by 2 to get x = b/2
that's your solution.
you can confirm by giving b any value and then making x that value divided by 2 and then solving the original equation.
i set b = 500 and x = 250.
original equation becomes 250/500 + 250/500 = 1 which becomes 500/500 = 1 which becomes 1 = 2, confirming the solution is correct.